OpenStudy (doc.brown):

In a population of test scores not known to be normally distributed, the mean is 50 and the standard deviation is 5. What proportion of the observations will fall in the interval from 40 to 60?

OpenStudy (doc.brown):

Does this have to do with the Z score? \[p(\left|Z\right|\ge k)\le\frac{1}{k^2}\]Or whatever this means?

OpenStudy (doc.brown):

@Loser66 Any idea what this mess is?

OpenStudy (loser66):

No idea, :)

OpenStudy (doc.brown):

I know how you feel.